# UPS: Efficiently Building Foundation Models for PDE Solving via Cross-Modal Adaptation

### v2

@@ 6 Conclusion and Future Work
We identify several future directions based on our work. First, we can validate our method on a broader range of PDEs with higher-order temporal derivatives or three-dimensional domains. Meanwhile, to seek a truly general foundation model for PDE, we aim to extend the types of tasks that UPS can solve. Currently, UPS is only applicable to forward prediction. It is important to study inverse problems of parameter estimation for different PDEs as well.
7 Broader Impact
@@ Acknowledgement
Appendix
Appendix A Explanation of Figure. 1
We use the results in Table 1 to generate the figure. For the y-axis, we first find out the maximum and minimum nRMSEs achieved by all considered methods. Then, we normalize the error $e$ for each method by taking $\frac{e-\min}{\max-\min}$. The dots represent the average values of the normalized errors across 7 PDE families from PDEBench. The error bars represent standard deviation. The x-axis of the plot represents our estimation of computational costs: we compute the product of the number of GPUs used and the number of training steps. We then take $log_{10}$ to make the visualization clearer.
Appendix B Datasets
As mentioned in Section 5, we train our models using the datasets provided in the PDEBench [10]. The time-dependent PDE families considered by our models are: Burgers Equation (1D), Diffusion-Sportion (1D), Shallow-Water (2D), compressible Navier-Stokes (1D and 2D), incompressible Navier-Stokes (2D), and Diffusion-Reaction (1D and 2D). For each $s\in S$, the number of points in the $n$-point discretization $W_{n}^{s}$ is $128$, i.e, $n=128$. For PDEs where the PDEbench-provided grid has more than $128$ points in each dimension, we sample $128$ equispaced points.
In this section, we provide few key properties and considerations for the PDEs used in this paper. The initial conditions $\bm{u}(0,x)$ for most of the datasets are sampled from a superposition of sinusoidal waves. The set of coefficients and number of trajectories used per PDE are reported in Appendix Table 5. For full details on the data generation process and the hyperparameters used to generate the PDE dataset, we refer the reader to Takamoto et al. [10].
B.1 Burgers Equation (1D)
@@
B.2 Diffusion-Sorption Equation (1D)
@@
where $x\in(0,1)$, $t\in(0,500]$, and $D=5\times 10^{-4}$. For more details on the initial conditions, boundary conditions and the function $R(u)$, we refer the reader to Takamoto et al. [10]. For our training, we use $4500$ trajectories for this PDE generated by varying the initial conditions.
B.3 Advection Equation (1D)
@@
where $x\in(0,1)$ and $t\in(0,2]$. Various examples in this dataset are generated by sampling multiple initial conditions from a super-position of sinusoidal waves as used in  Takamoto et al. [10].
B.4 Compressible Navier-Stokes (1D and 2D)
@@
B.5 Incompressible fluid Navier-Stokes (2D)
@@
For density $\rho$ and pressure $p$, the equations used to generate the data in Takamoto et al. [10] are as follows:
@@
where $\bm{f}$ is an external forcing function, and Dirichlet boundary conditions. Here $x\in[0,1]^{2}$ and the initial conditions $\bm{u}$ and the forcing term $\bm{f}$ are sampled from two-dimensional Gaussian random fields. Please refer to Takamoto et al. [10] for more details on the data generation process.
B.6 Reaction Diffusion (1D and 2D)
@@
B.7 Shallow-Water Equations (2D)
@@
$\begin{split}&\partial_{t}h+\partial_{x_{1}}hu_{1}+\partial_{x_{2}}hu_{2}=0,\\ &\partial_{t}hu_{1}+\partial_{x_{1}}\left(u_{1}^{2}h+\frac{1}{2}g_{r}h^{2}\right)=-g_{r}h\partial_{x_{1}}b,\\ &\partial_{t}hu_{2}+\partial_{x_{2}}\left(u_{2}^{2}h+\frac{1}{2}g_{r}h^{2}\right)=-g_{r}h\partial_{x_{2}}b,\end{split}$ (13)
@@ B.8 Summary
Appendix C Training Details
C.1 Hyperparameters
@@
  • Stage 1 epoch: 20
  • Stage 2 epoch: 100
We use the CoNLL-2003 dataset [48] as the reference dataset for alignment in stage 1.
C.2 Compute
We run all of our experiments on a single NVIDIA A6000. Below are the detailed model size, per epoch training time (in seconds), and total training time (in hours). Note that we train the models for 100 epochs.
@@ C.2 Compute
Appendix D Detailed Experiment Results
D.1 2D-Only UPS
Table 7: Training UPS with all of the 2D datasets in PDEBench and compare with MPP and DPOT. Note that beyond these PDEBench datasets, MPP is also pretrained on PDEArena [11] and DPOT is pretrained on PDEArena [11] as well as CFDBench [49]. Baseline results taken from Hao et al. [14]. ‘–’ means that the result is not available.
@@ D.2 Few-Shot Adaptation
D.3 Ablation on Longer Sequence Length
@@
Appendix E Visualization
E.1 Burgers Equation
E.2 Diffusion-Sorption
E.3 1D Navier Stokes
@@ E.2 Diffusion-Sorption
E.4 Shallow Water
E.5 2D Navier Stokes

### v4

@@ 6 Conclusion and Future Work
We identify several future directions based on our work. First, we can validate our method on a broader range of PDEs with higher-order temporal derivatives or three-dimensional domains. Meanwhile, to seek a truly general foundation model for PDE, we aim to extend the types of tasks that UPS can solve. Currently, UPS is only applicable to forward prediction. It is important to study inverse problems of parameter estimation for different PDEs as well. For an impact statement, see Appendix A.5.
@@ Acknowledgement
Appendix A Appendix
A.1 Datasets
As mentioned in Section 5, we train our models using the datasets provided in the PDEBench (Takamoto et al., 2022). The time-dependent PDE families considered by our models are: Burgers Equation (1D), Diffusion-Sportion (1D), Shallow-Water (2D), compressible Navier-Stokes (1D and 2D), incompressible Navier-Stokes (2D), and Diffusion-Reaction (1D and 2D). For each $s\in S$, the number of points in the $n$-point discretization $W_{n}^{s}$ is $128$, i.e, $n=128$. For PDEs where the PDEbench-provided grid has more than $128$ points in each dimension, we sample $128$ equispaced points.
In this section, we provide few key properties and considerations for the PDEs used in this paper. The initial conditions $\bm{u}(0,x)$ for most of the datasets are sampled from a superposition of sinusoidal waves. The set of coefficients and number of trajectories used per PDE are reported in Appendix Table 5. For full details on the data generation process and the hyperparameters used to generate the PDE dataset, we refer the reader to Takamoto et al. (2022).
A.1.1 Burgers Equation (1D)
@@
A.1.2 Diffusion-Sorption Equation (1D)
@@
where $x\in(0,1)$, $t\in(0,500]$, and $D=5\times 10^{-4}$. For more details on the initial conditions, boundary conditions and the function $R(u)$, we refer the reader to Takamoto et al. (2022). For our training, we use $4500$ trajectories for this PDE generated by varying the initial conditions.
A.1.3 Advection Equation (1D)
@@
where $x\in(0,1)$ and $t\in(0,2]$. Various examples in this dataset are generated by sampling multiple initial conditions from a super-position of sinusoidal waves as used in  Takamoto et al. (2022).
A.1.4 Compressible Navier-Stokes (1D and 2D)
@@
A.1.5 Incompressible Fluid Navier-Stokes (2D)
@@
For density $\rho$ and pressure $p$, the equations used to generate the data in Takamoto et al. (2022) are as follows:
@@
where $\bm{f}$ is an external forcing function, and Dirichlet boundary conditions. Here $x\in[0,1]^{2}$ and the initial conditions $\bm{u}$ and the forcing term $\bm{f}$ are sampled from two-dimensional Gaussian random fields. Please refer to Takamoto et al. (2022) for more details on the data generation process.
A.1.6 Reaction Diffusion (1D and 2D)
@@
A.1.7 Shallow-Water Equations (2D)
@@
$\begin{split}&\partial_{t}h+\partial_{x_{1}}hu_{1}+\partial_{x_{2}}hu_{2}=0,\\ &\partial_{t}hu_{1}+\partial_{x_{1}}\left(u_{1}^{2}h+\frac{1}{2}g_{r}h^{2}\right)+\partial_{x_{2}}u_{1}u_{2}h=-g_{r}h\partial_{x_{1}}b,\\ &\partial_{t}hu_{2}+\partial_{x_{2}}\left(u_{2}^{2}h+\frac{1}{2}g_{r}h^{2}\right)+\partial_{x_{1}}u_{1}u_{2}h=-g_{r}h\partial_{x_{2}}b,\end{split}$ (13)
@@ A.1.8 Summary
A.2 Experiment Details
A.2.1 Training Hyperparameters
@@
  • Training epoch: 20 for stage 1, 100 for stage 2
We use the CoNLL-2003 dataset (Sang & Meulder, 2003) as the reference dataset for alignment in stage 1.
A.2.2 Efficiency Analysis
We run all of our experiments on a single NVIDIA A6000. Table 6 show the detailed model size, per epoch training time (in seconds), and total training time (in hours) for different network configurations. Note that we train the models for 100 epochs.
@@ A.2.2 Efficiency Analysis
We reported additional metrics such as FLOPs and the time required for predicting a single step for a PDE instance in Table 7, assuming the input data is 2D with 4 channels and resolution 128. We mainly compared with unified models that have similar model sizes. Compared to these existing work, UPS has lower FLOPs and shorter inference time. This shows that our model is ideal for deployment in practical environments where both computational efficiency and speed are critical.
Table 7: Efficiency comparison for unified neural operators.
UPS-B MPP-B DPOT-M
Num Params 149M 116M 122M
Per Forward Pass FLOPs (G) 72.66 102.12 75.44
Single Step Inference Time (ms) 1.77 2.34 1.88
A.3 Detailed Experiment Results
A.3.1 2D-Only UPS
Table 8: Training UPS with all of the 2D datasets in PDEBench and compare with MPP and DPOT. Note that beyond these PDEBench datasets, MPP is also pretrained on PDEArena (Gupta & Brandstetter, 2022) and DPOT is pretrained on PDEArena (Gupta & Brandstetter, 2022) as well as CFDBench (Yining et al., 2023). Baseline results taken from Hao et al. (2024). ‘-’ means that the result is not available.
@@ A.3.2 Few-Shot Adaptation
A.3.3 Ablation on Longer Sequence Length
@@
A.3.4 Long-Horizon Prediction
As stated in Section 3, our method mainly focuses on predicting the next step from the current step, i.e., $\hat{\bm{u}}^{s}_{t+1}(\bm{x})={\mathcal{G}}_{\theta}(\bm{u}^{s}_{t}(\bm{x}))$. However, we are also interested in the prediction capacity of our method over a longer period of time. Thus, we study an additional setting that predicts $\hat{\bm{u}}^{s}_{t+10}(\bm{x})$. We show non-autogressive evaluation results since otherwise we will only have very few time steps for each test PDE trajectory. The table below shows that UPS is still effective for long-horizon prediction compared to baselines like FNO. Even though the prediction interval is longer, the error rates only slightly increase possible because we use non-autoregressive evaluation, so the errors do not accumulate.
1D 1D 1D 1D 2D 2D 2D
$\Delta t=1$ 0.0027 0.0399 0.0009 0.0056 0.0016 0.0153 0.0931
$\Delta t=10$ 0.0034 0.04 0.0011 0.0074 0.0026 0.0189 0.134
A.4 Visualization
A.4.1 Burgers Equation
A.4.2 Diffusion-Sorption
A.4.3 1D Navier Stokes
@@ A.4.2 Diffusion-Sorption
A.4.4 Shallow Water
A.4.5 2D Navier Stokes